OpenStudy (anonymous):

Evaluate 10x²y - 1 for x = 3 and y = 2

OpenStudy (dumbcow):

179

OpenStudy (anonymous):

its 45

OpenStudy (anonymous):

10*3^2 *2-1=10*9*2-1=180-1=179

OpenStudy (anonymous):

That's not an answer choice. A. 2/9 B. 45 C. -45

OpenStudy (anonymous):

it's y^-1 sorry guys.

OpenStudy (anonymous):

10*3^2/2= 90/2 = 45